Joe has x CDs, Penny has y CDs and Jahal has z CDs.
(a) If (x+y)(x+z)=x^2+10x+24, find the values of y and z.
(b) How many CDs should Joe have if the total number of CDs belonging to the three friends is fifteen?

Answer by ankor@dixie-net.com(22740)

Joe has x CDs, Penny has y CDs and Jahal has z CDs.
;
(a) If (x+y)(x+z) = x^2+10x+24, find the values of y and z.
factor x^2 + 10x + 24
(x+6)(x+4)
Therefore we can say
y = 6 and z = 4 (or vice-versa)
:
(b) How many CDs should Joe have if the total number of CDs belonging to the three friends is fifteen?
y & z have a total of 10 cds therefore Joe (x) has 5